Please help ASAP.

A new treatment has been developed to alleviate the symptoms of Hansen’s disease (once known as Leprosy), which is prevalent on Island Z. The treatment does not eliminate all symptoms but substantially reduces their severity. All of those with the disease had been advised to take a one-time treatment that costs $5,000 (Treatment A). However a new treatment has come on the market. It is also taken one time, but like many drugs with a small market and patent protection it is very expensive, costing $100,000 (Treatment B). Hansen’s disease is painful and disfiguring. Surveys of those on the island who have taken Treatment A reveal that they on average are indifferent between 10 years of life with their symptoms and 5 years of life symptom - free. In a pilot program for Treatment B, it was found that, due to reduced severity of the symptoms, those who had taken the new treatment were indifferent to 10 years with the disease and 7 years of life symptom-free. The average age of onset of Hansen’s disease is 40. Life expectancy overall on the island is 75, but at age 40 life expectancy is 77 and at age 55 life expectancy is 79.

1, Calculate the ICER for the Treatment B in Quality Adjusted Life years (QALYs) for those newly diagnosed with the disease at the average age of onset.

2. Calculate the ICER for Treatment B for those suffering from the disease at the average age for this population of 55.

3. To justify Treatment B through cost-benefit analysis for the person who experiences onset of the disease at the average age of 40, what would be the necessary minimum valuation for a quality adjusted life year?

4. How would the ICERs change if it was found that half of those who take the treatments do not respond at all?

5. How would the ICERs change if life expectancy increased by 10 percent (e.g. 75 X 1.1)?
Answer #1

1. The ICER for the treatment B in Quality Adjusted Life Years (QALY),

ICER = (C1 - C0)/(E1 - E0); C1, E1 are cost and effect of the intervening group and C0,E0 are cost and effect of the control care group.

ICER = ( 1,00,000 - 5000)/(55 - 40) = 95,000/15 = 6333.34

2. The ICER for Treatment B for those suffering from the disease at the average age of the population of 55,

ICER = (1,00,000 - 5000)/(79 - 77) = 95,000/2 = 47,500.

4. If half the population who takes the treatment dies not respond at all, then the effect of the treatment will decrease. Thus the ICER will increase its value.

5. If the expectancy increased by 10%, then the below value the effect will increase, and thus the ICER will increase in value.